A study reports that 36% of companies in Country A have three or more female board directors. Suppose you select a random sample of 100 respondents. Complete parts​ (a) through​ (c) below.

a) What is the probability that the sample will have between 35% and 38% of companies in Country A that have three or more female board​ directors?

The probability is .2460

The probability is 80% that the sample percentage of Country A companies having three or more female board directors will be contained within what symmetrical limits of the population​percentage?

The probability is 80% that the sample percentage will be contained above what ​% and below what ​%.
Answer #1

Ans:

a)

z(0.35)=(0.35-0.36)/sqrt(0.36*(1-0.36)/100)=-0.208

z(0.38)=(0.38-0.36)/sqrt(0.36*(1-0.36)/100)=0.417

P(-0.208<z<0.417)=P(z<0.417)-P(z<-0.208)

  =0.6615-0.4175=0.2440

b)

lower limit=0.36-1.282*sqrt(0.36*(1-0.36)/100)=0.2985 or 29.85%

upper limit=0.36+1.282*sqrt(0.36*(1-0.36)/100)=0.4215 or 42.15%